The ratio in which the point (10/7,33/7)divides the line segment joining the points (1,3) and (2 ,7) is p:q then p+q is (p and q are coprime)? I have solved it but there are coming two eqn which is equal so i can`t find p and q!

Grade:12th pass

1 Answers

Pinki Shah
6 years ago
2p+q/p+q=10/7. 7(2p+q)=10(p+q) 14p+7q=10p+10q. 4p=3q. . P/q=3/4. Hence p+q = 3+4=7. Asdghjkiuyreefcvbjutredcvbjjirdccbjutdfcbjfdvnkitddcbjurdcvnhrd